OpenStudy (anonymous):

if n=5 find the value of n+n+n this is confusing :/

OpenStudy (diyadiya):

15

OpenStudy (anonymous):

5+5+5=15

OpenStudy (anonymous):

n + n +n 5 + 5+5

OpenStudy (anonymous):

equal 15

jimthompson5910 (jim_thompson5910):

n+n+n = 5+5+5 = 15 or n+n+n = 3n = 3(5) = 15